OpenStudy (anonymous):

Jonathan uploaded some original songs and also some pictures. Alissa uploaded four times as many songs, but only uploaded twice as many pictures. Jonathan uploaded 11 items while Alissa uploaded 24. How many songs did Jonathan upload?

OpenStudy (anonymous):

x + 3y = 9 3x − 3y = −13 I don't get this one either

OpenStudy (wolfe8):

This is a simultaneous equation thingy(forgot the name). First you'd want to give symbols for the songs and pictures. I will do x for songs Jonathan uploaded, y for pictures Jonathan uploaded. With the info you have, Alissa uploaded 4 times songs Jon did, so that's 4x. Twice songs so 2y. In the end you are told that Jon uploaded 11 items(songs AND pictures). So x + y = 11 For Alissa, 4x + 2y = 24 So now you have 2 equations: x + y = 11 4x + 2y = 24 Can you do it from here? Solve for one of the variables first then the second one. There are the elimination method and substitution method.

OpenStudy (wolfe8):

Oh I'm sorry. Since the question wants hoe many songs Jon uploaded, you only need to solve for x.